the speed of a boat in still water is 28 mph. the boat traveled 70 mi down a river in the same amount of time it took to travel 42 mi up the river. find the rate of the river's current.
Answer by jorel1380(3719)

70/28+c=42/28-c
1960-70c=1176+42c
112c=784
c=7
The river's current is 7 mph..